Quadratic Functions Word Problem
2. Find two numbers whose difference is 14 and whose product is a minimum. use "Let Statement". Thank you!!!

Answer by Alan3354(69443)

Find two numbers whose difference is 14 and whose product is a minimum.
Prod = x*(x-14)
P = x^2 - 14x
dP/dx = 2x - 14 = 0
x = 7
2nd number = -7